# AI-Based Stroke Rehabilitation Domiciliary Assessment System with ST-GCN Attention

### v1

@@ IV-A1 Dataset
We conduct experiments on data scored on a 50-point scale based on the Likert method. The datasets include both whole-body and upper-body rehabilitation exercises. whole-body tasks reflect global motor coordination, while upper-body tasks emphasize fine motor control of the arms and hands. conducting experiments on the two datasets, we evaluate the model performance in assessing different rehabilitation exercises. a) KIMORE Kinematic assessment of Movement for remote monitoring of physical Rehabilitation, called KIMORE[35] is a dataset consisting RGB-D videos and score annotations for five whole-body exercises. The Subject groups(GPP) are divided according to the presence or absence of pain or disability. The non-pain group(CG) includes motions performed by 12 professionals include physiotherapist, as well as 32 non-professionals. The GPP group consists of 34 subjects with conditions such as Parkinson’s disease, back pain, and stroke. Score annotations for each exercise were provided by a professional rehabilitation therapist. b) NRC(ours) This dataset consists of upper-limb rehabilitation movement data collected using RGB-D cameras, IMU sensors, and dedicated software. It includes 15 exercises comprising 10 activities of daily living (ADL) and 5 range of motion (ROM) tasks in Table. I . The ADL set emphasizes upper-body and finger movements relevant to daily life. We further categorized the exercises into three groups: UNI (Unimodal Hand Using), BIA (Bimodal and Asymmetric Hand Using), and BIS (Bimodal and Symmetric Hand Using). These categories enable comparative analysis of the model performance in assessment score mappings across different types of movements. All motion data include score annotations provided by professional rehabilitation therapists. For this purpose, we developed a new set of 10 dedicated evaluation items.
We designed an end-to-end system for rehabilitation recognition and assessment using deep learning, with the proposed model corresponding to the rehabilitation assessment module. For this task, the model was trained on data from six non-disabled (ND) subjects and eleven stroke patients. The effectiveness of rehabilitation depends on the time of onset[38]. This factor represents a main requirement for the goal of our system, which is to enable patients to perform rehabilitation exercises independently at home. Accordingly, patient recruitment focused primarily on subacute stroke patients with hemiplegia within six months of onset. The detailed conditions for the recruited stroke participants are as follows:
@@
In total, 1,142 movement–score pairs were used, segmented by exercise. The dataset was split into train, validation and test sets at approximately 8:1:1 ratio. A summary of the datasets is provided in Table I.
@@ IV-A2 Preprocessing
a) Pose estimation The raw data consists of RGB-D video sequences, which requires an additional step of skeleton keypoint estimation. For this purpose, we employed MMPose[39], an open-source human pose estimation library developed by the OpenMMLab team. This framework provides a unified interface for researchers and developers to train, infer, and evaluate a wide variety of pose estimation models. In our work, we performed inference using the HRNet-DarkPose model [40, 41], which was trained on the COCO-WholeBody dataset[31], to estimate human poses. The extracted skeleton structure is presented in Fig. 1. b) Uniform Frame Sequences Standardizing frame sequences into a uniform format is a critical preprocessing step for deep learning models. Although individual actions may differ in temporal characteristics (e.g., length, speed), consistency across the dataset is required. Inevitably, this process involves partial loss of motion information. To mitigate this, we applied a frame drop preprocessing strategy. As illustrated in Fig. 4., the original sequence was divided into groups at intervals of $N$ frames, and a frame was sampled from each group and concatenated sequentially. This approach preserves the original temporal tendency. It also allows frame shifting for random sampling within groups, which enhances data variability and supports augmentation. Using this method, all sequences were standardized to a length of 288 frames.
Figure 4: Uniform Frame Sequence Method.
@@ IV-B Training
The implementation and training were conducted on an Intel i9-10920X CPU and an NVIDIA GeForce RTX 4090 GPU. All model frameworks were implemented in PyTorch on Ubuntu 22.04.05 LTS. The hyperparameters were set as follows: 200 epochs, batch size of 64, and AdamW optimizer with a learning rate of 0.003. For performance evaluation, Mean Absolute Devian (MAD), Root Mean Square Error (RMSE), and Mean Absolute Percentage Error (MAPE) were employed. These metrics quantify the discrepancy between ground truth and model predictions in terms of mean, root square, and percentage error, respectively. The equations are as follows:
@@ V Result
TABLE IV: COMPARISON OF ASSESSMENT DEEP LEARNING MODEL RESULTS ON KIMORE USING RMS, MAPE AND MAD METRICS (LOWER VALUES INDICATE BETTER PERFORMNACE)
@@ V-A Overview
We carefully distinguished between baseline models in our paper. Deep learning research for rehabilitation assessment remains a task with significant potential, yet reliable prior studies are still restrictive. Accordingly, we employed existing human-task deep learning models for performance comparison. The models listed in Table IV under the Recognition Model–FC layer category correspond to this approach. References [13, 45, 43, 44, 46] were designed for recognition tasks. For assessment, their last softmax layers were replaced with fully connected layers. This is possible since both assessment and recognition rely on extracting qualitative information from human action data. Since human behavior is a high-level construct involving multiple modalities, several studies have explored multimodal deep learning rather than relying on a single modality. Kuang et al. proposed a hierarchical approach that separately learned position and orientation and then integrated them for joint representation.
@@ V-A1 KIMORE Dataset Result
we compared RAST-G@ with baseline models using three metrics, (6), (7), (8). The results in Table IV, The model improved accuracy and maintained consistent performance across all five actions, indicating the model’s potential applicability not only to full-body tasks but also to diverse rehabilitation movements and assessment domains. Among the three evaluation metrics, MAD showed modest improvement over previous models, whereas RMSE and MAPE demonstrated substantially larger gains. This is because of the metric definitions. MAD averages absolute errors and is less sensitive to large deviations, whereas RMSE and MAPE capture the influence of large errors and their proportion relative to overall predictions.
The results of Ex4 also provide an important indicator. Relative Works have reported particularly low errors for this action, which can be explained by its nature. This Exercise is pelvis rotations on the transverse plane while keeping other body parts fixed, where rotational variation provides better representation than translational displacement. In a Cartesian representation, depth (z-axis)serves as the primary feature, whereas in a quaternion representation, the motion can be more effectively expressed through roll and pitch axes. Thus, the method of Kuang et al. was well suited for this action, showed very small errors.
These result finds that our model is not specialized for a single action but instead maintains robust performance across a diverse set of movements. Its stability becomes more pronounced as the number of predicted actions increases and as inter-action variability grows. Consequently, RAST-G@ is advantageous for stroke rehabilitation assessment, where patient-specific movements differ and long-term monitoring is required.
@@
Metric RAST-G@ Kuang et al. Deb et al.
@@ V-A2 NRC(Ours) Dataset Result
The NRC dataset collected for this study includes activities of daily living (ADL) and range of motion (ROM) exercises. ADL tasks consist of relatively dense and complex movements and are generally more difficult compared to other actions. They are further constrained to the upper extremities, concentrating motion information in the arms and fingers. For baseline, we compared our method with graph-based deep learning approaches proposed by Kuang et al.,[42] and Deb et al.,[26]. Kuang et al. employed a hierarchical strategy that separately learned position and orientation, later integrating the two for joint representation. Orientation was represented by four-dimensional quaternions $(x,y,z,w)$ which show the directional state of each joint[47]. Accordingly, it was necessary to generate appropriate orientation data from the NRC dataset prior to experimentation. Algorithm 1. describes this procedure. Given $J$ skeletons, a skeleton sequence seqseq, and an edge tuple array $E$, the function BUILD_PARENTS takes $J$ and $E$ as input and returns an index array $P$ indicating the parent joint of each joint $j$. The function QUATERNION_SEQUENCE then takes $seq,P$, and the previous skeleton frame prevprev from, i1i-1 as input, and returns a sequence of joint-wise quaternions. A reference REST pose is computed to serve as the baseline for rotation estimation. For each joint $j$ the parent–child connection is used to calculate the rotation from the baseline pose. A reference vector $up_{0}$ is introduced for computing the roll angle around the $y$-axis, ensuring orthogonality of the $x,y,z$ rotation axes. This process is repeated for each frame $T$ producing quaternion sequences. Finally, the angular differences between the reference pose and the current pose are computed to form matrices $D_{t,j}$ , which are transformed into $Q\in\mathbb{R}^{T\times J\times 4}$.
When trained on the NRC dataset with both position and orientation features, RAST-G@ achieved substantial improvements over baseline models. This result means the characteristics of the NRC dataset, unlike repetitive tasks, includes goal-directed movements accompanied by fine finger control. Emphasizing temporal modeling led to superior performance, as summarized in Table. V. Whereas other baseline models exhibited error increases of up to fivefold when transferred to the NRC dataset, RAST-G@ maintained stable accuracy, thereby demonstrating robustness. Moreover, the model consistently outperformed baselines in error measures that account for outliers (e.g., RMSE) and relative error magnitude (e.g., MAPE). Notably, MAPE even decreased further, underscoring the model’s potential for rehabilitation domains characterized by high input variability.
Fig. 5. presents the ground-truth scores provided by therapists and the predictions of RAST-G@ about NRC dataset. A longitudinal study was conducted once per month over four months, during which action data and therapist scores for our assessment were collected. The comparison shows that the model not only learned the overall trend of score but also provided good quality score predictions.
@@ V-B Ablation Study
In Table. VI., The frame dimension is a critical factor for representing action sequences, and the strategy used to construct uniform input sequences directly affects the model’s perception of motion. Frame dropping, which reduces sequence length while preserving the overall temporal flow, is one of the core components of RAST-G@. Without frame dropping, the model yielded the lowest MAPE, although MAD and RMSE did not differ substantially from the proposed method. This indicates that preprocessing without considering full action sequences can distort temporal information and result in larger relative errors.
The temporal Attention aimed to learn fine-grained rehabilitation movements by leveraging a deeper neural network architecture. without the method showed the most significant performance degradation, underscoring the importance of temporal information for aligning rehabilitation motions with qualitative assessment. This emphasizes that effective modeling of temporal dynamics is essential for deep learning systems to interpret rehabilitation tasks and, more broadly, human action data.
@@
frame Drop 0.260 0.326 0.379
Temporal Attention 0.277 0.340 0.480
@@
27 Keypoint represents a more exploratory component of the ablation study rather than direct evidence for our claims. Because the definition of “important” landmarks in the human body can be subjective, it was necessary to evaluate whether our custom skeleton design for the upper extremities, derived from COCO WholeBody, was appropriate for the collected data. Moreover, increasing the number of keypoints does not necessarily yield better data representation. Prior studies have noted that not all joints are equally informative and that redundant joints can introduce noise [48, 49]. Attention mechanisms have been applied to emphasize key joints within the overall context, showing that selective focus is more effective than uniform treatment. In the case of hands, which are densely articulated, excessive keypoints may even degrade performance. For instance, [50] employed graph reduction in skeleton-based sign language recognition, reporting a drop in accuracy to 63.69 when redundant nodes were removed. Since our dataset includes ADL tasks involving fine-grained hand movements, we compared model performance across skeleton configurations using our proposed 27 keypoint upper-limb representation. While larger numbers of keypoints theoretically provide more information, the results show only marginal gains for MAD and RMSE, while MAPE worsened. This implies that although absolute error magnitudes may decrease, relative accuracy compared to action scale diminishes, reflecting errors at the global contextual level. These findings suggest that for high-dimensional sequential motion data, reducing complexity while maintaining structural consistency is more effective than indiscriminately increasing keypoints.
Unlike other approaches[16, 17, 18, 19, 20, 22] that score actions by comparing input with reference movements, our method focuses on learning the mapping between action patterns and scores from training data. As shown in the Stroke-only ablation study (Table VI), training on Stroke only data produced error rates comparable to the Proposal(Table V). Because both training and testing were conducted on Stroke data, direct error metrics such as MAD observed even better results. In contrast, RMSE and MAPE exhibited higher errors than those obtained with mixed ND-Stroke data, indicating that ND sequences are more effective for capturing movement flow, whereas Stroke patient data tend to include more abrupt movements.
Conclusionally the RAST-G@ with optimized frame construction (via frame dropping), temporal attention, and a 25 keypoint skeleton, preserved low MAD and RMSE while significantly reducing MAPE. This demonstrates that the model not only controls absolute error but also achieves higher relative accuracy with respect to actual motion magnitude, which is critical for rehabilitation assessment and broader human action understanding.
@@ V-C Feedback for User
In proposing a domiciliary rehabilitation system, we also considered the need for user-friendly feedback. Model outputs must be interpretable not only to clinicians but also to patients and elderly users who may have limited familiarity with artificial intelligence, computer systems, or medical terminology. Building on the framework, our system includes a mobile application that provides comprehensive user feedback(Fig. 6). The feedback is organized into two main components: period feedback and discrete feedback.
Period feedback summarizes performance trends over time. A representative score, calculated as the average, is reported for each month to visualize changes in rehabilitation progress. The duration can be adjusted according to user preference, and performance graphs are available for overall scores as well as for specific categories such as OVERALL, UNI, BIA, and BIS. Since difficulty varies across movement categories, this enables more fine-grained evaluation. Such longitudinal feedback can assist therapists in designing and adjusting treatment plans for individual patients.
Discrete feedback records the outcome of each exercise, including skeleton heatmap visualizations, task-specific scores, and—optionally—therapist comments. This provides patients with intuitive and interpretable results, while also allowing clinicians to document and track qualitative aspects of performance. Fig. 7. illustrates the video feedback interface of our rehabilitation system. The results, derived from the assessment module including RAST-G@, highlight the most influential joints and edges in the motion sequence $m$ at frame $t_{1}$. These are represented as heatmaps overlaid on the skeleton sequence, where color intensity and node size indicate their relative contribution to the assessment. The input tensor $X\in\mathbb{R}^{N\times C\times T\times V\times 1}$ passes through the ST-GCN blocks to capture skeletal structure (spatial direction) along with limited temporal information, which forms the core learning representation while preserving the skeleton configuration. The resulting feature representation, $X^{L}\in\mathbb{R}^{{(N\times M)}\times 256\times{(T/4)}\times V}$ is used to generate skeleton heatmaps for visual feedback. For interpretability, the model score was rescaled to a 0–100 scale. In beta testing with physiotherapists, the system was positively received, particularly for its ability to visualize postural alignment and joint utilization. Beyond enhancing user-friendliness, the visual feedback also provides an important means to observe and analyze how the model attends to different body regions during learning.

### v2

@@ IV-A1 Dataset
We conduct experiments on datasets scored on a 50-point Likert-based scale. The evaluated datasets include both whole-body and upper-limb rehabilitation exercises. Whole-body tasks reflect global motor coordination, whereas upper-limb tasks emphasize fine motor control of the arms and hands. By conducting experiments on these two datasets, we assess the model performance across different types of rehabilitation actions. a) KIMORE Kinematic Assessment of Movement for Remote Monitoring of Physical Rehabilitation (KIMORE)[6] is a dataset comprising RGB-D videos and score annotations for five whole-body exercises. Participants are grouped according to the presence or absence of pain or disability. The control group (CG) includes movements performed by 12 professionals, including physiotherapists, and 32 non-professionals. The group with pain and postural disorders(GPP) consists of 34 participants with conditions such as Parkinson’s disease, back pain, stroke. For each exercise, scores were provided by a professional rehabilitation therapist.
b) NRC(ours) This dataset consists of upper-limb rehabilitation movement data collected using RGB-D cameras, IMU sensors, and dedicated software.[18] It includes 15 exercises comprising 10 ADL and 5 ROM tasks, as listed in Table II. The ADL set emphasizes upper-limb and finger movements relevant to daily life. We further categorized the exercises into three groups: UNI (unimanual hand use), BIA (bimanual asymmetric hand use), and BIS (bimanual symmetric hand use). These categories enable comparative analysis of model performance in mapping diverse action types to assessment scores. All action samples in both datasets are annotated with scores provided by professional rehabilitation therapists. For NRC, we developed a new set of 10 dedicated assessment items. The assessment score is derived from these items.
We designed an end-to-end deep learning system for rehabilitation action recognition and assessment, where the proposed model serves as the assessment module. For this task, the model was trained on data collected from 6 ND participants and 11 stroke patients. The effectiveness of rehabilitation depends on the time since stroke onset[17] and repetitive practice. This consideration motivates the primary goal of our system: enabling patients to perform rehabilitation exercises independently at home. In particular, participant recruitment focused on subacute stroke with hemiplegia within six months of stroke onset. The detailed inclusion criteria for the recruited stroke participants are as follows:
@@
In total, 1,142 movement–score pairs were used, segmented per exercise. The dataset was split into train, validation, and test sets at approximately 8:1:1 ratio. A summary of the datasets is provided in Table I.
@@ IV-A2 Preprocessing
a) Pose estimation The raw data consist of RGB-D video sequences, which require an additional skeleton keypoint estimation step. For this purpose, we employed MMPose, an open-source human pose estimation library developed by the OpenMMLab team.[32] This framework provides a unified interface for training, inference, and evaluation across a wide range of pose estimation models. In our work, we ran inference using the HRNet-DarkPose model[50, 40], trained on the COCO-WholeBody dataset[45], to estimate human poses. The extracted skeleton representation is presented in Fig. 1. b) Uniform Frame Sequences Standardizing frame sequences into a uniform format is a critical preprocessing step for deep learning models. Although individual actions may differ in temporal characteristics such as duration and speed, consistency across the dataset is necessary. Inevitably, this process involves partial information loss. To mitigate this, we applied a frame-dropping preprocessing strategy. As illustrated in Fig. 4, the original sequence was divided into non-overlapping groups of $N$ frames. One frame was then sampled from each group and concatenated in temporal order. This approach preserves the overall temporal trend of the action. It also enables random frame selection within each group, which increases data diversity and facilitates augmentation. Using this strategy, all sequences were standardized to a length of 288 frames. The raw data distribution is reported in another study of ours.[18]
Figure 4: Uniform frame sequence method by frame dropping.
@@ IV-B Training
The implementation and training were conducted on an Intel i9-10920X CPU and an NVIDIA GeForce RTX 4090 GPU. All models were implemented in PyTorch and trained on Ubuntu 22.04.5 LTS. The hyperparameters were set as follows: 200 epochs, a batch size of 64, and the AdamW optimizer with a learning rate of 0.003. For performance evaluation, Mean Absolute Deviation (MAD), Root Mean Square Error (RMSE), and Mean Absolute Percentage Error (MAPE) were employed. The equations are as follows:
@@
Each metric is computed over the entire test set and reported as an average error between ground-truth scores and model predictions.
@@ V Result
TABLE IV: COMPARISON OF ASSESSMENT DEEP LEARNING MODEL RESULTS ON KIMORE USING RMSE, MAPE AND MAD METRICS (LOWER VALUES INDICATE BETTER PERFORMNACE)
@@ V-A Overview
We carefully selected and categorized baseline models for comparison. Although deep learning–based rehabilitation assessment remains a promising research area, the number of well-validated prior studies is still limited. Accordingly, we adopted established human action recognition models as baselines. The models listed in Table IV under the Recognition Model–FC layer category correspond to this setting. The methods[46, 51, 38, 26, 49] were originally designed for recognition tasks; for assessment, we replaced their softmax based classification heads with fully connected layers to perform score regression. This adaptation is feasible because both recognition and assessment rely on extracting discriminative features from human action data. Because human behavior is a high-level data that can be captured through multiple modalities, several studies have explored multimodal deep learning rather than relying on a single modality. For example, Kuang et al.[22] proposed a hierarchical approach that separately learns position and orientation and then integrates them to form a joint representation.
@@ V-A1 KIMORE Dataset Result
We compared our model, RAST-G@, with baseline models using the three evaluation metrics defined in (6), (7), (8). As reported in Table IV, the proposed model achieved lower errors and maintained consistent performance across all five exercises. These results suggest that the model is applicable not only to whole-body tasks but also to a broader range of rehabilitation movements and assessment settings. Among the three metrics, MAD showed only modest improvement over prior models, whereas RMSE and MAPE exhibited substantially larger gains. This difference can be attributed to how the metrics summarize prediction errors. MAD averages absolute deviations and is less sensitive to large errors, while RMSE and MAPE more strongly reflect the impact of large deviations and their relative magnitude. Regarding Ex4, the results can be explained by its motion characteristics. Prior work has reported particularly low errors for this exercise, which is consistent with its motion characteristics. Ex4 mainly involves pelvic rotation in the transverse plane while other body parts remain relatively fixed. In this setting, rotational variation may be more informative than translational displacement. Under a Cartesian representation, z-axis can dominate the motion signal, whereas quaternion-based representations can capture rotational components more directly. Consequently, the method of Kuang et al. is well suited for Ex4 and yields very small errors.
Overall, these results indicate that our model is not specialized for a single exercise but instead maintains robust performance across a diverse set of actions. This robustness becomes more evident as the number of evaluated exercises increases and inter-exercise variability grows. Accordingly, RAST-G@ is advantageous for stroke rehabilitation assessment, where patient-specific movement patterns vary substantially and long-term monitoring is required.
@@
Metric RAST-G@ Kuang et al.[22] Deb et al.[10]
@@ V-A2 NRC(Ours) Dataset Result
The NRC dataset collected in this study includes both ADL and ROM exercises. ADL tasks involve relatively complex and fine-grained movements and may be more challenging than other actions, depending on stroke severity. Moreover, they are limited to the upper extremities, such that motion information are concentrated in the arms and fingers. For baseline comparisons, we evaluated our method against graph-based deep learning approaches proposed by Kuang et al.[22] and Deb et al.[10]. Kuang et al. employed a hierarchical strategy that learns position and orientation separately and then integrates them into a joint representation. Orientation was represented using four-dimensional quaternions $(x,y,z,w)$, which encode the directional state of each joint.[36] Accordingly, we converted the Cartesian skeleton coordinates into a quaternion-based representation prior to experimentation. Given a skeleton sequence seqseq and the skeletal connectivity defined by parent–child joint relations, we first computed a reference REST pose to serve as the baseline for rotation estimation. For each frame, joint orientations were estimated relative to this baseline by comparing each parent–child bone direction with the corresponding direction in the REST pose. To obtain a stable and consistent rotation parameterization, a reference vector $up_{0}$ was introduced to determine the roll component around the $y$-axis and to enforce orthogonality among the $x,y$, and $z$-rotation axes. Repeating this process over all frames produces a sequence of joint-wise quaternions. The resulting orientation sequence is represented as $Q\in\mathbb{R}^{T\times J\times 4}$, where each element corresponds to a quaternion describing the orientation of joint $j$ at time $t$. When trained using both position and orientation features, RAST-G@ showed substantial improvements over the baseline models on the NRC dataset. NRC contains target actions that require fine finger control and exhibit temporal structure, which allows the proposed temporal modeling strategy to be particularly effective. As summarized in Table V, while other baselines exhibited up to a fivefold increase in error when transferred to NRC, RAST-G@ maintained stable performance, demonstrating robustness under dataset shift. Moreover, RAST-G@ consistently outperformed the baselines on error measures that emphasize large deviations and relative error magnitude. Notably, MAPE was lower than the baselines, indicating robustness across subjects and task types. Fig. 5 presents the ground-truth scores provided by therapists and the corresponding predictions of RAST-G@ on the NRC dataset. A longitudinal study was conducted monthly over four months, during which movement data and therapist-assigned scores were collected for assessment. The comparison indicates that the model captures the overall trend in scores and predicts accurate score.
@@ V-B Ablation Study
The ablation study in Table VI provides a comparison of the effect of each method. The frame standardization strategy used to unify input can affect the model’s predictions. Frame dropping reduces sequence length while preserving the overall temporal progression. Without frame dropping, MAPE exhibited a increase, whereas MAD and RMSE remained unchanged. This indicates inadequate temporal preservation during standardization distorts action features and increases prediction error.
The TA block was introduced to capture temporal dependencies in rehabilitation movements. Without the TA block caused the largest performance degradation, highlighting the importance of temporal information when mapping motions to qualitative assessment scores. Overall, this indicates that temporal modeling is essential for understanding rehabilitation tasks and more broadly, human actions.
@@
w/o Frame Drop 0.260 0.326 0.379
w/o Temporal Attention(TA) 0.277 0.340 0.480
Keypoints: 27 0.261 0.318 0.427
@@
The Keypoint 27 setting was introduced as a comparison factor for optimizing data representation, rather than for validating the proposed method itself. As shown in Fig. 1, we derived skeleton from the COCO WholeBody configuration. However, our target movements differ from those in the benchmark datasets used for pose estimation. We therefore evaluated whether the skeleton is appropriate for our data. Prior work has emphasized the value of selecting informative landmarks. [28, 47] introduced attention mechanisms to focus on important keypoints in full-body motion, and related efforts in sign-language modeling address dense hand regions. [16] used a skeleton graph with a hand-focused structure reporting a drop in accuracy to 63.69 when apply graph node reduction. Since our dataset includes ADL tasks involving fine-grained hand motions, we evaluated a comparable configuration and compared the resulting performance. However, MAPE decreased, indicating improved relative accuracy. This suggests that, for human action tasks requiring holistic motion understanding, the choice of skeleton representation should prioritize consistency over increased dimensionality. For sequential action data with substantial temporal and spatial complexity, reducing redundant information while preserving overall structure can be more effective.
Unlike reference-based approaches[14, 39, 52, 1, 35, 23] that assign scores by comparing an input sequence to an ideal movement template, our method learns a direct mapping from action patterns to assessment scores using supervised training data. As shown in the stroke-only ablation results, training on stroke data alone achieved performance comparable to the proposed setting. Because both training and testing were performed on stroke sequences, absolute-error metrics such as MAD remained similar and in some cases slightly improved. In contrast, RMSE and MAPE were higher than in the ND–stroke setting, suggesting that including ND sequences helps the model capture smoother movement flow, whereas stroke sequences often contain more abrupt temporal variations. Overall, RAST-G@ with optimized frame construction, the TA module, and a 25-keypoint skeleton maintained low MAD and RMSE while reducing MAPE. The model not only limits absolute error but also improves accuracy with respect to motion magnitude, which is important for rehabilitation assessment and broader human action understanding.
@@ V-C Feedback for User
In designing a domiciliary rehabilitation system, we also considered the need for user-friendly feedback. Model outputs should be interpretable not only to clinicians but also to patients and older adults who may have limited familiarity with artificial intelligence, computer systems, or medical terminology. Based on this requirement, we developed a mobile application that provides structured feedback to users as shown in Fig. 6. The feedback is organized into two components: period feedback and discrete feedback.
V-C1 Period feedback
Period feedback summarizes performance trends over time. For each month, a representative score computed as an average is reported to visualize changes in rehabilitation progress. The time window can be adjusted according to user preference, and performance graphs are available for both overall scores and the categories OVERALL, UNI, BIA, and BIS. Because task difficulty differs across categories, this view enables more fine-grained monitoring. Such longitudinal feedback can also support therapists in designing and adjusting individualized treatment plans.
V-C2 Discrete feedback
Discrete feedback records the outcome of each exercise by providing skeleton heatmaps, task-specific scores, and, when available, therapist comments. This feedback offers patients an intuitive summary of performance and enables clinicians to document qualitative observations over time. Fig. 7 presents the vizualization feedback interface of the proposed system. The visual feedback is generated from the assessment module, including RAST-G@, and highlights the most influential joints and edges in the motion sequence $m$ at frame $t_{1}$. These regions are displayed as heatmaps overlaid on the skeleton sequence, where color intensity and node size indicate their contribution to the assessment. The input tensor $X\in\mathbb{R}^{N\times C\times T\times V\times 1}$ is processed by ST-GCN blocks to capture skeletal structure and temporal context while preserving the skeleton configuration. The resulting feature representation $X^{L}\in\mathbb{R}^{{(N\times M)}\times 256\times{(T/4)}\times V}$ is then used to generate skeleton heatmaps for visual feedback. For clarity, the score is rescaled to a 0–100 range.
In beta testing with physiotherapists, the interface was well received, particularly for visualizing postural alignment and joint utilization. Beyond improving usability, this feedback enables analysis of how the model attends to different body regions during inference, supporting interpretability and informing future model analysis and refinement.